# Principles for Open Data Curation: A Case Study with the New York City 311 Service Request Data

### v1

@@ 1 Introduction
In the early 21st century, the open data movement began to take shape, driven by the fundamental belief that freely accessible data can transform both societies and governments. This movement champions the principles of transparency, innovation, and public engagement. A landmark in this journey was the launch of the United States’ Data.gov portal in 2009 (General Services Administration,, 2024), a pioneering platform in making government data widely accessible. Shortly afterwards, the European Union followed suit, unveiling its Open Data Portal (European Union Plublications Office,, 2024) in 2012, further cementing the movement’s global reach. Furthermore, the World Bank’s Open Data initiative, initiated in 2010, stands out as a comprehensive repository for global development data, available at World Bank Open Data (World Bank,, 2024). These initiatives represent significant strides in democratizing data, removing barriers that once kept valuable information regarding government performance in silos. Their collective impact extends beyond mere data sharing to fostering a culture of openness that benefits individuals, communities, governments, and economies worldwide (Barns,, 2016; Wang and Lo,, 2016). Despite these global efforts, the effective use of open data hinges on addressing critical challenges in data curation, such as ensuring quality, consistency, and usability.
@@
The City of New York (NYC) has emerged as a leader in the open data movement, marked by the enactment of the Open Data Law in 2012 (Zuiderwijk and Janssen,, 2014). This landmark legislation led to the creation of the NYC Open Data Portal (NYC Open Data Team,, 2024), which today hosts an impressive array of 2,700 datasets across 80 different city agencies. This wealth of data serves as a powerful tool for researchers and policymakers, significantly enhancing local government transparency. Popular datasets include information on restaurant health inspection violations, car crashes, high school and college enrollment statistics, jail inmate charges, and the location of City-wide free Internet access points. These datasets have been applied in civil life in various ways, such as mapping car crashes involving pedestrians and visualizing high school and college enrollment trends. Furthermore, they have enabled significant research across multiple domains, including health (Cantor et al.,, 2018; Shankar et al.,, 2021), urban development (Neves et al.,, 2020), and transportation (Gerte et al.,, 2019), aiding in the understanding and addressing of complex urban challenges. For example, NYC’s 311 data has been used to optimize resource allocation in urban planning and improve emergency response times, showcasing its practical relevance in real-time decision-making.
@@
However, the open data presents substantial curation challenges. Data curation, the process of organizing, maintaining, and ensuring the quality of datasets, plays a crucial role in maximizing the utility of open data. Proper curation ensures that datasets remain consistent, accurate, and useful for diverse applications. For example, well-curated data is essential for machine learning systems, which require high-quality data to produce reliable insights (Polyzotis et al.,, 2019; Jain et al.,, 2020). A critical component of data curation is data cleaning, which involves identifying and rectifying inconsistencies, errors, and inaccuracies in datasets. While open data initiatives have made vast amounts of data available, ensuring the reliability and utility of such data hinges on rigorous data cleaning processes. Efficient usage of software tools can significantly streamline this aspect of curation (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018). Poor curation may result in issues such as missing data, formatting errors, or inconsistent values, leading to biased or inaccurate outcomes (Geiger et al.,, 2020). This is particularly critical in domains where machine learning is applied to sensitive tasks, such as public health or policy (Rahm and Do,, 2000).
@@
Research into data curation has explored these challenges in-depth. Among the earliest discussions, Witt et al., (2009) focused on developing data curation profiles tailored to specific contexts, setting a precedent for targeted data management strategies. Addressing broader challenges in data sharing and management, Borgman, (2012) highlighted the complexities of research data distribution, emphasizing the need for robust strategies. This is complemented by Hart et al., (2016), who outlined essential principles for effective data management, particularly emphasizing meticulous curation practices. The utility of curated open data is vividly illustrated in public health and global challenges, where Cantor et al., (2018) demonstrated the utility of curated data in evaluating community health determinants, and Shankar et al., (2021) observed its critical role during the COVID-19 pandemic in managing collective responses. Despite these advances, there remains a notable gap in providing actionable frameworks for managing large-scale municipal datasets.
@@
The paper is organized as follows. Section 2 provides an overview of the history of the NYC 311 SR system and presents a summary of SR counts over a 2-year period from 2022 to 2023. Section 3 delves into specific data cleansing challenges affecting data quality and curation efficiency, including structural problems, adherence to the data dictionary, and the presence of missing, blank, or N/A entries. This section also investigates data field compliance with reference or acceptable values, highlights logical inconsistencies, and examines some concerning patterns in the data. We explore the balance between precision and accuracy and identify duplicate or redundant fields, along with observations on the Data Dictionary (NYC Open Data Team,, 2023). Section 4 offers practical recommendations for mitigating or resolving these issues, while Section 5 encapsulates key insights and discusses the broader implications of our findings.
@@ 2 NYC 311 Service Request (SR) Data
The impact of NYC 311 data extends beyond operational efficiency; it has become instrumental in shaping City governance and community engagement. Open data not only ensures governmental transparency but also empowers civic developers, the general public, and policy-makers (Minkoff,, 2016; O’Brien et al.,, 2017; Kontokosta and Hong,, 2021). The data has been pivotal in providing advice on shelters during emergencies, handling inquiries during the COVID-19 pandemic, enforcing standards between landlords and tenants, reallocating taxi routes based on analyses by the Taxi and Limousine Commission (TLC), and improving responsiveness across City Agencies. It also supports a wide range of studies in resource allocation and emergency response strategies (Zha and Veloso,, 2014; Raj et al.,, 2021), social equity in service provision (White and Trump,, 2018; Kontokosta and Hong,, 2021), and urban challenges such as noise pollution (Dove et al.,, 2022) and street flooding (Agonafir et al.,, 2022).
@@ 3 Data Cleansing Issues
Data cleansing refers to the process of identifying and rectifying errors, inconsistencies, and inaccuracies within datasets to ensure they are of high quality and reliable for analysis (Maletic and Marcus,, 2005; Hosseinzadeh et al.,, 2023). The process typically involves removing duplicate records, handling missing or incomplete data, correcting mislabeled or inaccurate entries, and standardizing data formats (e.g., Cody,, 2017; Van der Loo and De Jonge,, 2018). In the context of open data, cleansing is especially important as open datasets often come from diverse, often uncoordinated sources, leading to variations in data quality, completeness, and consistency. Without thorough data cleansing, the utility of open data can be limited and perhaps untrustworthy, affecting its reliability for research, policy making, and innovation. The main purpose of cleansing open data is to ensure that it is accurate, consistent, and usable a cross multiple analytical platforms and by various stakeholders with a myriad of purposes. Data cleansing improves the trustworthiness of the data and enables more accurate analysis, better decision making, and the improved integration of data into machine learning models or other systems.
@@ 3.3 Validating Data for Acceptable Values
(b) Post-closed resolution_action_update_date(s) ¿30 days
@@ 3.6 Reducing File Size
One promising alternative to reducing data storage for this large 311 SR dataset (3.7GB) would be employment of the cross-platform Apache Arrow package. Apache Arrow (arrow.apache.org) is an innovative, cross-language development platform designed for in-memory data. It defines a standard for representing columnar data, enabling efficient analytics and real-time data processing. As such, it offers dramatic improvements in file storage size, as well as efficient analysis and cross-platform support for R, Python, and Julia programming languages. Remarkable storage size reduction (within several orders of magnitude) is possible (Bates and Yan,, 2024).
@@ 5 Discussion
Our study highlights the critical role of robust data curation in ensuring the reliability and utility of open datasets, as demonstrated through the examination of the NYC 311 data. As these datasets are applied across diverse fields such as public services and academic research, inconsistencies, missing values, and formatting errors can significantly undermine the insights derived from them. Trustworthy machine learning systems, which rely heavily on high quality data, are especially vulnerable to such issues. Errors in the underlying data, such as biases or inconsistencies, can compromise both the accuracy and fairness of predictive models, impacting real-world decisions in urban planning and policy (Rahm and Do,, 2000; Geiger et al.,, 2020).
While the findings here are specific to NYC 311 data, the principles proposed have broader applicability. Datasets in healthcare, transportation, and environmental monitoring can similarly benefit from these curation strategies, enabling improved outcomes across domains. This research further emphasizes the necessity of harmonizing data across City agencies, particularly resolving discrepancies in field names, formats, usage, and definitions. Consistency in data is essential for trustworthy machine learning, where inconsistencies across sources can lead to distorted outcomes. This challenge is even more pronounced in datasets that span long time periods, as changes in agency structures or reporting standards which can introduce subtle biases and data distortions. Long-term data consistency is crucial for longitudinal studies and predictive modeling, as even minor data shifts can lead to significant deviations in model outcomes, as noted by Rahm and Do, (2000) andBorgman, (2012).
Finally, the intersection of data curation and machine learning for public policy applications opens new avenues for improving governance. High quality data enables machine learning models to better predict trends, allocate resources, and address service delivery issues, such as delays in responding to 311 complaints. The COVID-19 pandemic underscored the importance of real-time, trustworthy data in crisis management, where the accuracy and timeliness of data-driven insights were critical for public health responses. Without proper data curation, machine learning models used during the pandemic would have been compromised, affecting decisions on resource allocation and service provision (Worby and Chang,, 2020; Khemasuwan and Colt,, 2021). Future efforts could focus on automating data curation processes, particularly in real-time data pipelines, to ensure that the data used in machine learning models remains accurate, clean, and reliable (Chu et al.,, 2016; Hurbean et al.,, 2021).

### v2

@@ Abstract
Data cleansing,
Data science,
NYC Open Data,
Smart city,
Transparency,
doi:
$0$
keywords:
volume: 0issue: 0articletype: research-article
@@ 1 Introduction
In the early 21st century, the open data movement began to take shape, driven by the fundamental belief that freely accessible data can transform both societies and governments. This movement champions the principles of transparency, innovation, and public engagement. A landmark in this journey was the launch of the United States’ Data.gov portal in 2009 (General Services Administration, 2024), a pioneering platform in making government data widely accessible. Shortly afterwards, the European Union followed suit, unveiling its Open Data Portal (European Union Plublications Office, 2024) in 2012, further cementing the movement’s global reach. Furthermore, the World Bank’s Open Data initiative, initiated in 2010, stands out as a comprehensive repository for global development data, available at World Bank Open Data (World Bank, 2024). These initiatives represent significant strides in democratizing data, removing barriers that once kept valuable information regarding government performance in silos. Their collective impact extends beyond mere data sharing to fostering a culture of openness that benefits individuals, communities, governments, and economies worldwide (Barns, 2016; Wang and Lo, 2016). Despite these global efforts, the effective use of open data hinges on addressing critical challenges in data curation, such as ensuring quality, consistency, and usability.
@@
The City of New York (NYC) has emerged as a leader in the open data movement, marked by the enactment of the Open Data Law in 2012 (Zuiderwijk and Janssen, 2014). This landmark legislation led to the creation of the NYC Open Data Portal (NYC Open Data Team, 2024), which today hosts an impressive array of 2,700 datasets across 80 different city agencies. This wealth of data serves as a powerful tool for researchers and policymakers, significantly enhancing local government transparency. Popular datasets include information on restaurant health inspection violations, car crashes, high school and college enrollment statistics, jail inmate charges, and the location of City-wide free Internet access points. These datasets have been applied in civil life in various ways, such as mapping car crashes involving pedestrians and visualizing high school and college enrollment trends. Furthermore, they have enabled significant research across multiple domains, including health (Cantor et al., 2018; Shankar et al., 2021), urban development (Neves et al., 2020), and transportation (Gerte et al., 2019), aiding in the understanding and addressing of complex urban challenges. For example, NYC’s 311 data has been used to optimize resource allocation in urban planning and improve emergency response times, showcasing its practical relevance in real-time decision-making.
@@
However, the open data presents substantial curation challenges. Data curation, the process of organizing, maintaining, and ensuring the quality of datasets, plays a crucial role in maximizing the utility of open data. Proper curation ensures that datasets remain consistent, accurate, and useful for diverse applications. For example, well-curated data is essential for machine learning systems, which require high-quality data to produce reliable insights (Polyzotis et al., 2019; Jain et al., 2020). A critical component of data curation is data cleaning, which involves identifying and rectifying inconsistencies, errors, and inaccuracies in datasets. While open data initiatives have made vast amounts of data available, ensuring the reliability and utility of such data hinges on rigorous data cleaning processes. Efficient usage of software tools can significantly streamline this aspect of curation (e.g., Cody, 2017; Van der Loo and De Jonge, 2018). Poor curation may result in issues such as missing data, formatting errors, or inconsistent values, leading to biased or inaccurate outcomes (Geiger et al., 2020). This is particularly critical in domains where machine learning is applied to sensitive tasks, such as public health or policy (Rahm and Do, 2000).
@@
Research into data curation has explored these challenges in-depth. Among the earliest discussions, Witt et al. (2009) focused on developing data curation profiles tailored to specific contexts, setting a precedent for targeted data management strategies. Addressing broader challenges in data sharing and management, Borgman (2012) highlighted the complexities of research data distribution, emphasizing the need for robust strategies. This is complemented by Hart et al. (2016), who outlined essential principles for effective data management, particularly emphasizing meticulous curation practices. The utility of curated open data is vividly illustrated in public health and global challenges, where Cantor et al. (2018) demonstrated the utility of curated data in evaluating community health determinants, and Shankar et al. (2021) observed its critical role during the COVID-19 pandemic in managing collective responses. Despite these advances, there remains a notable gap in providing actionable frameworks for managing large-scale municipal datasets.
@@
The paper is organized as follows. Section 2 provides an overview of the history of the NYC 311 SR system and presents a summary of SR counts over a 2-year period from 2022 to 2023. Section 3 delves into specific data cleansing challenges affecting data quality and curation efficiency, including structural problems, adherence to the data dictionary, and the presence of missing, blank, or N/A entries. This section also investigates data field compliance with reference or acceptable values, highlights logical inconsistencies, and examines some concerning patterns in the data. We explore the balance between precision and accuracy and identify duplicate or redundant fields, along with observations on the Data Dictionary (NYC Open Data Team, 2023). Section 4 offers practical recommendations for mitigating or resolving these issues, while Section 5 encapsulates key insights and discusses the broader implications of our findings.
@@ 2 NYC 311 Service Request (SR) Data
The impact of NYC 311 data extends beyond operational efficiency; it has become instrumental in shaping City governance and community engagement. Open data not only ensures governmental transparency but also empowers civic developers, the general public, and policy-makers (Minkoff, 2016; O’Brien et al., 2017; Kontokosta and Hong, 2021). The data has been pivotal in providing advice on shelters during emergencies, handling inquiries during the COVID-19 pandemic, enforcing standards between landlords and tenants, reallocating taxi routes based on analyses by the Taxi and Limousine Commission (TLC), and improving responsiveness across City Agencies. It also supports a wide range of studies in resource allocation and emergency response strategies (Zha and Veloso, 2014; Raj et al., 2021), social equity in service provision (White and Trump, 2018; Kontokosta and Hong, 2021), and urban challenges such as noise pollution (Dove et al., 2022) and street flooding (Agonafir et al., 2022).
@@ 3 Data Cleansing Issues
Data cleansing refers to the process of identifying and rectifying errors, inconsistencies, and inaccuracies within datasets to ensure they are of high quality and reliable for analysis (Maletic and Marcus, 2005; Hosseinzadeh et al., 2023). The process typically involves removing duplicate records, handling missing or incomplete data, correcting mislabeled or inaccurate entries, and standardizing data formats (e.g., Cody, 2017; Van der Loo and De Jonge, 2018). In the context of open data, cleansing is especially important as open datasets often come from diverse, often uncoordinated sources, leading to variations in data quality, completeness, and consistency. Without thorough data cleansing, the utility of open data can be limited and perhaps untrustworthy, affecting its reliability for research, policy making, and innovation. The main purpose of cleansing open data is to ensure that it is accurate, consistent, and usable a cross multiple analytical platforms and by various stakeholders with a myriad of purposes. Data cleansing improves the trustworthiness of the data and enables more accurate analysis, better decision making, and the improved integration of data into machine learning models or other systems.
@@ 3.3 Validating Data for Acceptable Values
(b) Post-closed resolution_action_update_date(s) >30 days
@@ 3.6 Reducing File Size
One promising alternative to reducing data storage for this large 311 SR dataset (3.7GB) would be employment of the cross-platform Apache Arrow package. Apache Arrow (arrow.apache.org) is an innovative, cross-language development platform designed for in-memory data. It defines a standard for representing columnar data, enabling efficient analytics and real-time data processing. As such, it offers dramatic improvements in file storage size, as well as efficient analysis and cross-platform support for R, Python, and Julia programming languages. Remarkable storage size reduction (within several orders of magnitude) is possible (Bates and Yan, 2024).
@@ 5 Discussion
Our study highlights the critical role of robust data curation in ensuring the reliability and utility of open datasets, as demonstrated through the examination of the NYC 311 data. As these datasets are applied across diverse fields such as public services and academic research, inconsistencies, missing values, and formatting errors can significantly undermine the insights derived from them. Trustworthy machine learning systems, which rely heavily on high quality data, are especially vulnerable to such issues. Errors in the underlying data, such as biases or inconsistencies, can compromise both the accuracy and fairness of predictive models, impacting real-world decisions in urban planning and policy (Rahm and Do, 2000; Geiger et al., 2020).
While the findings here are specific to NYC 311 data, the principles proposed have broader applicability. Datasets in healthcare, transportation, and environmental monitoring can similarly benefit from these curation strategies, enabling improved outcomes across domains. This research further emphasizes the necessity of harmonizing data across City agencies, particularly resolving discrepancies in field names, formats, usage, and definitions. Consistency in data is essential for trustworthy machine learning, where inconsistencies across sources can lead to distorted outcomes. This challenge is even more pronounced in datasets that span long time periods, as changes in agency structures or reporting standards which can introduce subtle biases and data distortions. Long-term data consistency is crucial for longitudinal studies and predictive modeling, as even minor data shifts can lead to significant deviations in model outcomes, as noted by Rahm and Do (2000) andBorgman (2012).
Finally, the intersection of data curation and machine learning for public policy applications opens new avenues for improving governance. High quality data enables machine learning models to better predict trends, allocate resources, and address service delivery issues, such as delays in responding to 311 complaints. The COVID-19 pandemic underscored the importance of real-time, trustworthy data in crisis management, where the accuracy and timeliness of data-driven insights were critical for public health responses. Without proper data curation, machine learning models used during the pandemic would have been compromised, affecting decisions on resource allocation and service provision (Worby and Chang, 2020; Khemasuwan and Colt, 2021). Future efforts could focus on automating data curation processes, particularly in real-time data pipelines, to ensure that the data used in machine learning models remains accurate, clean, and reliable (Chu et al., 2016; Hurbean et al., 2021).